Reports until 17:23, Monday 16 December 2024
GQuEST Output Filter Cavities
Torrey Cullen - posted 17:23, Monday 16 December 2024 (12053)
OFC1 Cavity Realignment

[Jeff, Torrey]

Over the last week I have been struggling to get OFC1 realigned. The 1550 path can be locked on a very clean 0,0 mode. The 775 light on the other hand caused a few issues, the main problem being it was resonant but badly clipping on the piezo as it exited the cavity. The 775 light looked like it was hitting the piezo mirror at the top left of the mirror. We then locked the 1550 light and used the SWIR camera to look at the 775 output coupler. This showed alot of scattered light, but no solid beam. After breaking the cavity this scatter disappeared completely, suggesting the resonant path in the cavity hits the piezo upon exiting. We took off the 1550 input coupler to confirm this and it was hitting the same top left portion of the piezo mirror. 

With this information we have decided to adjust the inter cavity mirror alignment. The piezo mirror has a significant amount of down tilt since it was first contacted and re-set. Our plan tomorrow is to align the 775 input, but first take apart the piezo mirror and reassemble so that it sets in a different alignment.